7.58 Liquid nitrogen is an extremely cold liquid (-196 °C) used in chemistry magic shows to shrink balloons (and do other
Answer #1

Assuming pressure to be constant  Volume is directly proportional to its absolute temperature

V7

Hence V1/T1=V2/T2

V1 = 2L ; T1 = 26 + 273 = 299 K

T2 = Temperature of liquid nitrogen = -196 + 273 = 77 K
All temperatures must be in kelvin
2/299=V2 / 77

V2=2/299 x 77
=0.52 litres

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION